OpenStudy (anonymous):

Check my work: (7+x)(x^2+6x-8) Combine like terms to -1+7x+x^2 then what?

Nnesha (nnesha):

how did you get 7x ?? first step is to foil

Nnesha (nnesha):

distribute 2nd parentheses by first one

OpenStudy (anonymous):

6x+x is how I got 7x.

OpenStudy (anonymous):

=(7+x)(x2+6x+−8) =(7)(x2)+(7)(6x)+(7)(−8)+(x)(x2)+(x)(6x)+(x)(−8) =7x2+42x−56+x3+6x2−8x =x3+13x2+34x−56

OpenStudy (anonymous):

Okay so foil each?

Nnesha (nnesha):

 yep first multiply 2nd parentheses by 7